by TEI TEI

06/14/05 – Phasor turned in a new change order for Fire Alarm additional work at 250. The Change Order cost including all the additional inspections is $967.82. TE will approve and send to Diana for signing. Please return a signed copy to both ThermalTech and Phasor when complete.

06/08/05 – TTC sent final Reupert paperwork to Diana for complete project closeout. Noise issued on floor is being address by Mark Hardin.

05/11/05 – Sound issues arose from the toilet room and accompanying space being too quiet. GBBN made a field visit to investigate and the consultants gave recommendations about white noise or another exhaust fan. The County is to advise on how to proceed from here.

05/10/05 – TE emailed GBBN to advise that toilet exhaust on this project is connected to the main building toilet exhaust. If there is a problem have Tom check the main fan and if that doesn’t work, TE will measure the airflow and check the existing conditions contingent on a future request by the County.

05/02/05 – The building passed it fire alarm inspection this morning at 6:00 a.m. TTC, Jack Lamping, Darnell, Jim McFarland and Darryl Meadows were present and witnessed the test and the paperwork was signed. The City Inspector had issues with the existing system but did not hold up or not pass the test but these are issues, he wishes the County would fix because the system is technically not installed properly. Item (1) any smoke or pull device on the first floor reports as “Smoke & Pull Stations”. The inspector strongly urges that these two actuation devices be separated and report separately when in alarm. Item (2) when in alarm and the “Alarm Silence” button is pushed it disengage all the horns AND strobes on the first floor – this is against code as ONLY the horns should be disabled and the strobes should continue to flash until the system is RESET (it was shown many times during our testing periods that they are not wired properly). Item (3) They questioned the “fail-secure” status of the doors upstairs and downstairs. When the FA system is in alarm all the door holds release as they should and all the security access is disabled so that a key is required to enter the space. I confirmed with TA and MS that this is the correct operation for your building. The people inside can always egress out as there are strike bars or non locking door handles from the inside to the outside so none of that is required to change. Item (4) The first floor system is a “System Sentry” type system which was in place when Simplex received the second contract to renovate the second floor space. They did this by daisy chaining all the existing 4-wire devices through a ZAM (Zone Actuation Module) which is why they all report as the same device back at the Simplex panel. TE would recommend putting in a project right now that upgrades that existing system to 100% Simplex (which is a fully addressable system 4010 model). We estimate about a $23,700 project (30 devices @ $9000, 3000LF @ $9000, 6 ZAM’s @ $3000, Programming @ $1500, Permit and Testing @ $1200). Please advise.

04/28/05 -TE spent most of the day at the building trying to get the Simplex system working in conjunction with the System Sensor system. We tried many custom programs but in the end nothing worked perfectly. TTC has 3 ways to try to get system fixed this morning and will meet Simplex back on site to repair the system. (1) write a custom program that actually works, (2) replace the 4 smoke detectors with authentic Simplex devices, (3) add a UL listed manual reset to reset the 1st and Bsmt floors. TTC will keep DES up to date as we go through the day.

04/27/05 – The project failed the initial Fire Alarm test this morning. Not from device malfunction but the actual FA panel would not reset after receiving an alarm from a smoke head. The system was pre-tested using pull stations and reset fine on Tuesday. Because of this the City did not issue an approval. Phasor, TE and the County contacted Simplex for an emergency call into a technicians to repair or reprogram the main FA panel. Jack talked to Simplex and will be back on site as soon as they arrive today to work on the system to assist in figuring out what is going on and wrong. He believes that he was able to get the technician to understand what was happening and hopes to resolve the issue today. The inspectors agreed to come back this week to approve.

04/26/05 – Phasor bought two strobes locally and will install them today and have everything ready for tomorrow morning. Jack said it was a scramble but he found them and is ready to install. They will be ready for tomorrow’s test. At 2:48 p.m. they reported that everything is set for final FA inspection on Wed at 6:00 a.m.

04/19/05 – Final fire alarm test is scheduled for 04/27/05 at 6:00 a.m. Tom will be on hand from the County for FA issues. TE will attend as Owner’s Rep. Phasor will be EC. Copp stated that all egress issues are resolved and will be working fine for our final test.

04/15/05 – TE contacted Phasor about final completion and inspection by email to make sure everyone is ready for next week.

04/06/05 – TE and Phasor and County are coordinating final inspection details for April 25th. TE waiting on contract interior information from County before confirming

04/05/05 – Project Meeting today. Set new schedule for completion. JB Schmidt scheduled to be complete by April 15. Phasor to be work on 4/15 and 4/18-4/20. Final Inspections to be tentatively scheduled for 04/25/05. TE to contact Phasor for FA inspection coordination with JB. TE to contact Copp about schedule for plumbing space and general building space and to work on resolving software issues. County ES/Facilities to contact Sheriff’s Dept about Panic button for receptionist in ES.

03/25/05 – Reupert delivered some of the final paperwork with no payapp today. TE will contact them about this.

03/15/05 – Portions of the project have been on hold awaiting the replacement of drywall and water sealant. TE received final paperwork and O&M manuals including balance reports from Reupert. The manuals are not bound or copied. TE will copy and make sets for County use and turn them in at the end of the project with final paperwork and payapp from HVAC contractor.

01/31/05 – TE attended last Thursday’s Const Mtg. HVAC is complete. EC is nearly complete. PC is complete with working fixtures. Hot water reached the LAV in under 60 seconds. All fixtures appeared to drain properly. Security is nearly complete but Copp has some final work to do.

01/26/05 – TE faxed one copy and mailed three copies of the Affidavit of Release of Liens to Reupert.

01/24/05 – Jamie of Reupert called TE and indicated that she was preparing final document packet but did not know where to find the Affidavit of Release of Liens forms. TE to mail the Affidavit of Release of Liens to Reupert.

01/18/05 – TTC sent Reupert their punchlist item (balance report) and asked for their O&M’s.

12/27/04 – TTC stopped by the building to coordinate lighting layout with Jack Lamping. Some of the HVAC piping is too close to the ceiling grid and the deep parabolics don’t fit into the grid and below the ductwork and 3-way heating piping and valves. We remedied this situation by moving some lights over a grid space or 2. TTC has a copy of the new locations for as-built purposes. TTC believes the new layout will provide sufficient lighting levels and be a workable layout for the building occupants. The only outstanding issue with the lighting at this point is that the light nearest the column near the restroom can either be swapped out with a shallower fixture like the ones slated for the storage room and the lens be replaced with a paracube or (2) 2×2 fixtures can be used instead of (1) 2×4 fixture. The former would cost the project little or no money, the ladder may cost $200 – $300 out of the remaining allowance. We should have an a final answer later today.

12/20/04 – Ken Reupert out of the office, Jamie will have both of the Affidavit of Original or Subcontractors sent to ThermalTech today or tomorrow (12/21/04).

12/09/04 – Reupert revised Pay Applications 1 & 2 received in the mail. Voice message left for Ken Reupert to disregard earlier message and that the Pay Application will be held for the “Affidavit of Original or Subcontractor” form to be submitted.

12/09/04 – Reupert Pay Applications 1 & 2 Voided due to $600.00 allowance shown on application as a Change Order. Voice message left for Ken Reupert explaining that the Applications were incorrect and will need to be revised.

12/03/04 – TTC talked to Reupert today. All the above ceiling ductwork is complete at this point. The 3-way valves are all piped in. The controls on on schedule at this point. The GC will contact the MC when the diffusers and final connections are ready. The MC has to attach some angles around the fire dampers to meet code and should take of that work next week.

11/22/04 – Reupert called today to discuss the 3-way valves ordered on this project via the submittal change. TE instructed them to pipe the valves in as 3-way valves and use up to $1250 of the allowance for this work and to make the necessary scheduling changes to make this happen. Per our submittals to the MC and CC we changed the 2-way valves to 3-way because this building has constant speed pumps and to prevent dead heading the heating hot water system will require 3-way valves in most installs.

11/18/04 – Reupert called to ask for a size change in the return grill. TE approved a 24 x 12 return opening above the ceiling for this project. TE to verify on site at today’s meeting all their request

10/27/04 – Reupert submitted a $600 allowance request to rework the piping at the bulkhead pursuant to our last construction meeting. TE approved the request from the $2500 mechanical allowance for this project, new balance $1900.

Change order request in approved for security change order and is the amount is not being subtracted from the allowance. TEI issued a revised sketch that corrected the access controlled door locations.

10/21/04 – TE contacted Reupert about missing today’s construction mtg, making the next meeting for sure, submitting their schedule of values, change order pricing for relocating piping and VAV box submittals. Everything is in the works. Reupert is going to send us the schedule, the VAV boxes are already in their shop (we now have to approve them (don’t worry they will be fine)), they will begin working on the change order request.

10/13/04 – AIA Change Order for $1,854.46 mailed to Phasor

10/11/04 – The Plumbing Permit Number is: # 2004P07058.

10/08/04 – TEI to prepare electrical bulletin #1 for 1854.36

10/07/04 – GBBN provided the permit number for the project: The permit number is as follows: #2004P05650

10/05/04 – TEI received devices submittal via email from Phasor.

10/05/04 – DES confirmed that TE sent Lisa the stamped drawings for permit. TE sent them on Thursday 10/07/04.

09/29/04 – Jeanette provided a dimensioned drawing for HamCo plumbing to use for this project – TE reviewed and approved for use, TE to pickup plumbing fixtures from Bo for approval, TE to provide stamped drawings for permit use to Lisa, TE to coordinate revised security (Copp to upgrade entire building including 2 new doors) and TE to write change order for this work to Phasor, Reupert was not at meeting and TTC to confirm that Reupert will be at next meeting, TE requested that GBBN reschedule Oct 20 mtg as it conflicts with CH Archive mtg.

09/20/04 – Ben has asked TTC to coordinate water closet locations with Bo Blank.

09/14/04 – Copp is providing a price to upgrade the whole 250 building, this has to be coordinated through TE and Phasor since our electrical contractor has security money in their bid to do this work.

09/08/04 – Bo advised to remove sanitary so that Reupert can begin construction.

09/08/04 – Pre Construction Mtg completed.

09/03/04 – DES advised that contracts are approved. GBBN will setup pre-construction meeting.

09/28/04 – Lighting submittal approved 9/22/04

10/01/04 – Request for change order pricing from Reupert for Fire Damper

10/02/04 – Reupert gave a verbal price of $375 for the addition of a fire damper